Method of Elimination Problem Solving
18. What happens when you solve the system 2x+3y=6 and 6x+9y=0 by elimination? Use a graph in your explanation.
Pleasee and thank you very much...

Answer by eperette(173)

2x + 3y = 6
6x + 9y = 0
Lets cancel x's
-3(2x + 3y = 6)
-6x -9y = -18
6x + 9y = 0
*****************
0 + 0 = -18
0 = -18
this is impossible, therefore there is no solution to this problem which means that the two lines representing this equations never intersect with each other....this only happens when two lines are parallel to each other (they have the same slope)
Note if you graph both equations you will see they are parallel and their slopes are the same (-2/3)
*****************